What will be the hydrogen ion concentration of a solution obtained by mixing 500 ml of 0.2 M acetic acid and 500 ml of 0.4 sodium acetate? (Ka CH3COOH) = 1.8 × 10−5) 


        (A)    1.8 × 10−5


        (B)    0.9 × 10−6


        (C)    9 × 10−6


        (D)    1.6 × 10−4


Grade:10

1 Answers

Navjyot Kalra
10 years ago

CH3COOH ? CH3COO− + H+

    K_a=[CH_3 COO^- ][H^+ ]/[CH_3 COOH]

    ? [H+] = K_a=[acid]/[salt]

        = 1.8 × 10−5 × 0.1/0.2 = 9 × 10−6